OpenStudy (kirbykirby):

Question about correlation! Show correlation = +/- 1 iff Y = aX+b

OpenStudy (kirbykirby):

Show \(|\rho (X,Y)|=1\) iff Y=aX+b, where \(\rho(X,Y)\) is the correlation coefficient of random variables X and Y, and a,b are constants with a \(\neq 0\). I am able to show \(\rho =1\) ,but unable that it can also be -1. \(Cov(X,Y)\) \(=Cov(X,aX+b)\) \(=E[X(aX+b)]-E(X)E(aX+b)\) \(=aE(X^2)+bE(X)-E(X)[aE(X)+b]\) \(=aE(X^2)+bE(X)-a[E(X)]^2-bE(X)\) \(=a[E(X^2)-[E(X)]^2]\) \(=aVar(X)\) Now, \(\rho(X,Y)\) \(= \frac{Cov(X,Y)}{\sqrt{Var(X)}\sqrt{Var(Y)}}\) \(= \frac{aVar(X)}{\sqrt{Var(X)}\sqrt{Var(aX+b)}}\) the numerator using the above result \(=\frac{aVar(X)}{\sqrt{Var(X)}\sqrt{a^2Var(X)}}\) \(=\frac{aVar(X)}{\sqrt{Var(X)}\big (a\sqrt{Var(X)}\big )}\) \(=\frac{aVar(X)}{aVar(X)} = 1\) ... I don't know to get it to be also either -1

OpenStudy (kirbykirby):

Do I just need to stick an absolute bar on both sides of the equation for \(\rho (X,Y)\) and get \(\rho (X,y) = |1| = \pm 1\)? It seems trivial to do so for some reason o_O.

OpenStudy (kirbykirby):

Actually I don't know what I wrote in my last reply... lol :(... |1| = 1 only lol

OpenStudy (anonymous):

if a is negative, it will be negative on the top, but the bottom is the square root of a^2, which is always positive.

OpenStudy (kirbykirby):

ohhh right!! \(\sqrt{a^2}=|a|\) I always forget that